Construct a frequency distribution and a frequency histogram for the Construct a frequency distribution of the data. Use the minimum data data set using the indicated number of classes. Describe any patterns.entry as the lower limit of the first class. Number of classes: 8 Data set: Finishing times (in seconds) of 20 Class Frequency male participants in a 5K race 리다□ □ 1719 1632 1573 1474 2300 1445 1250 1801 1597 1775 1955 1520 1824 1620 1304 1819 2020 2084 1580 1482 4、다□ □ 8.
Answer #1

Answer

Arranging the data in ascending or increasing order, we get

1250
1304
1445
1474
1482
1520
1573
1580
1597
1620
1632
1719
1775
1801
1819
1824
1955
2020
2084
2300

Now, we need to calculate the class width to make the required frequency table

We know that class width = (highest value - lowest value)/(number of classes)

highest value is 2300 and lowest value is 1250 and number of classes = 8

so, we get  

Class width = (2300-1250)/8 = 131.25 or 132(rounded to next whole number)

So, we can write classes as

1250 to 1250+132 =1250 to 1382

Similarly, we can keep on adding 132 till 8th row

so, it becomes

Class Frequency
1250 - 1382
1382 - 1514
1514 - 1646
1646 - 1778
1778 - 1910
1910 - 2042
2042 - 2174
2174 - 2306

Now, we have to count number of data values which falls within specific class.

So, the above table becomes

Class Frequency
1250 - 1382 2
1382 - 1514 3
1514 - 1646 6
1646 - 1778 2
1778 - 1910 3
1910 - 2042 2
2042 - 2174 1
2174 - 2306 1